# ViMMRC 2.0 - Enhancing Machine Reading Comprehension on Vietnamese Literature Text

### v2

@@ 3.4 Questions Analysis
1 | Wh-Questions. This type of questions is about What (cái gì), Who (ai, người nào), Which (vật gì, con gì), How (như thế nào), and How many (Bao nhiêu). The correct option is often matched directly from the reading passages. | Question: Bố Dũng đến trường học để làm điều gì? (English: What does Dung’s Father do when he goes to school with his son?) A. Để họp phụ huynh đầu năm (To join the parental meeting). B. Để đưa đồ cho Dũng (To give things to Dung). C. Để chào người thầy giáo cũ (To visit an old teacher). D. Để đưa Dũng đi học (To take Dung to his school). |
2 | Questions with listed options. This type of question gives a list of options and asks people to choose the correct one, all of the above, or none of the above. To choose the correct options, people need to read a bunch of paragraphs or even all the passages. | Question: Ý nghĩa của truyện Sơn Tinh, Thủy Tinh là gì? (English: What is the meaning of Son Tinh Thuy Tinh folktale?) A. Giải thích hiện tượng lũ lụt ở nước ta hằng năm (To explain the appearance of storm and flood) B. Thể hiện ước nguyện của con người trong việc chế ngự thiên nhiên. (To show the dream of controlling the disaster of humans). C. Ca ngợi công lao dựng nước của các vua Hùng. (To honor the feat of Hùng kings in the construction of the nation) D. Cả A, B và C đều đúng. (All of the choices are correct) |
3 | Yes/No questions. This question have only two options: Yes (Có / Đúng) or No (Không / Sai). | Question: Cốt truyện của Lặng lẽ Sa Pa là cốt truyện có tính kịch tính, xung đột. Đúng hay sai? (English: The story of Lang le Sa pa is a dramatic story, isn’t it?) A. Đúng (Yes). B. Sai (No). |
4 | Others | Question 1: Hãy sắp xếp các chi tiết dưới đây theo đúng thứ tự xuất hiện trong truyện Sơn Tinh, Thủy Tinh. (English: Let arrange these facts from the Son Tinh Thuy Tinh story by order of appearance) 1. Hùng Vương thứ mười tám nêu ra yêu cầu về lễ vật. (Hùng King XVIII gives the requirements of the marriage gift) 2. Sơn Tinh đem lễ vật đến trước và cưới được vợ. (Son Tinh comes first and married the bride) 3. Vua Hùng tổ chức kén rể cho Mị Nương. (Hùng King XVIII organizes a wedding for his daughter) 4. Sơn Tinh – Thủy Tinh đánh nhau ròng rã mấy tháng trời. (Son Tinh and Thuy Tinh keep fighting for months) A. (1) - (2) - (3) - (4). B. (1) - (3) - (2) - (4). C. (3) - (1) - (2) - (4). D. (1) - (3) - (4) - (2). Question 2: Điền từ chỉ màu sắc thích hợp vào khổ thơ sau: Em yêu màu … Đồng bằng, rừng núi Biển đầy cá tôm, Bầu trời cao vợi (Choose the word that describe the correct color to fill the blank: ”Em yêu màu … Đồng bằng, rừng núi Biển đầy cá tôm, Bầu trời cao vợi”). A. Xanh (Blue). B. Cam (Orange). C. Trắng (White). D. Đỏ (Red). |
@@
1 | Wh-Questions | 445 |
2 | Questions with listed options | 96 |
3 | Yes/No questions | 5 |
4 | Others | 18 |
@@ 6.3 Discussion
1 | Passage: passage#01 Question: Trong ba đứa cháu, ai là đứa cháu được ông nhận xét là còn thơ dại? (In three children, who is claimed as innocent?) A. Vân B. Việt C. Xuân D. Cả 3 bạn trên (All of those above) | A | D |
2 | Passage: passage#02 Question: Dựa vào diễn biến sự việc và tâm trạng nhân vật, nên chia đoạn trích thành mấy đoạn nhỏ là hợp lí? (Based on the happenings of the event and the character’s mood, how many parts should the excerpt be divided into?) A. 2 đoạn (Trước cảnh thề nguyền – Cảnh thề nguyền) (2 parts (Before the oath - The oath)) B. 3 đoạn (Kiều sang nhà Kim Trọng – Kim tỉnh giấc – Thề nguyền) (3 parts (Kieu goes to Kim Trong’s house - Kim wakes up - The oath)) C. 4 đoạn (cứ 5 dòng thơ một đoạn) (4 parts (1 every 5 lines)) D. A và B đều được (A and B are correct) | A | D |
3 | Passage: passage#03 Question: Dấu câu nào không dùng để tách thành phần câu được mở rộng trong bài văn này? (What punctuation is not used to separate the additional information from sentences in this text?) A. Dấu ngoặc đơn (brackets) B. Dấu hai chấm (colon) C. Dấu phẩy (comma) D. Dấu ngoặc đơn và dấu phẩy (brackets and comma) | B | D |
4 | Passage: passage#04 Question:: Hãy sắp xếp các chi tiết dưới đây theo đúng thứ tự xuất hiện trong truyện Sơn Tinh, Thủy Tinh. (Let arrange these facts from the Son Tinh Thuy Tinh story by order of appearance) 1. Hùng Vương thứ mười tám nêu ra yêu cầu về lễ vật. (Hùng King XVIII gives the requirements of the marriage gift) 2. Sơn Tinh đem lễ vật đến trước và cưới được vợ. (Son Tinh comes first and married the bride) 3. Vua Hùng tổ chức kén rể cho Mị Nương. (Hùng King XVIII organizes a wedding for his daughter) 4. Sơn Tinh – Thủy Tinh đánh nhau ròng rã mấy tháng trời. (Son Tinh and Thuy Tinh keep fighting for months) A. (1) - (2) - (3) - (4). B. (1) - (3) - (2) - (4). C. (3) - (1) - (2) - (4). D. (1) - (3) - (4) - (2). | C | D |

### v3

@@ 3.4 Questions Analysis
| 1 | Wh-Questions. This type of questions is about What (cái gì), Who (ai, người nào), Which (vật gì, con gì), How (như thế nào), and How many (Bao nhiêu). The correct option is often matched directly from the reading passages. | Question: Bố Dũng đến trường học để làm điều gì? (English: What does Dung’s Father do when he goes to school with his son?) A. Để họp phụ huynh đầu năm (To join the parental meeting). B. Để đưa đồ cho Dũng (To give things to Dung). C. Để chào người thầy giáo cũ (To visit an old teacher). D. Để đưa Dũng đi học (To take Dung to his school). |
| 2 | Questions with listed options. This type of question gives a list of options and asks people to choose the correct one, all of the above, or none of the above. To choose the correct options, people need to read a bunch of paragraphs or even all the passages. | Question: Ý nghĩa của truyện Sơn Tinh, Thủy Tinh là gì? (English: What is the meaning of Son Tinh Thuy Tinh folktale?) A. Giải thích hiện tượng lũ lụt ở nước ta hằng năm (To explain the appearance of storm and flood) B. Thể hiện ước nguyện của con người trong việc chế ngự thiên nhiên. (To show the dream of controlling the disaster of humans). C. Ca ngợi công lao dựng nước của các vua Hùng. (To honor the feat of Hùng kings in the construction of the nation) D. Cả A, B và C đều đúng. (All of the choices are correct) |
| 3 | Yes/No questions. This question have only two options: Yes (Có / Đúng) or No (Không / Sai). | Question: Cốt truyện của Lặng lẽ Sa Pa là cốt truyện có tính kịch tính, xung đột. Đúng hay sai? (English: The story of Lang le Sa pa is a dramatic story, isn’t it?) A. Đúng (Yes). B. Sai (No). |
| 4 | Others | Question 1: Hãy sắp xếp các chi tiết dưới đây theo đúng thứ tự xuất hiện trong truyện Sơn Tinh, Thủy Tinh. (English: Let arrange these facts from the Son Tinh Thuy Tinh story by order of appearance) 1. Hùng Vương thứ mười tám nêu ra yêu cầu về lễ vật. (Hùng King XVIII gives the requirements of the marriage gift) 2. Sơn Tinh đem lễ vật đến trước và cưới được vợ. (Son Tinh comes first and married the bride) 3. Vua Hùng tổ chức kén rể cho Mị Nương. (Hùng King XVIII organizes a wedding for his daughter) 4. Sơn Tinh – Thủy Tinh đánh nhau ròng rã mấy tháng trời. (Son Tinh and Thuy Tinh keep fighting for months) A. (1) - (2) - (3) - (4). B. (1) - (3) - (2) - (4). C. (3) - (1) - (2) - (4). D. (1) - (3) - (4) - (2). Question 2: Điền từ chỉ màu sắc thích hợp vào khổ thơ sau: Em yêu màu … Đồng bằng, rừng núi Biển đầy cá tôm, Bầu trời cao vợi (Choose the word that describe the correct color to fill the blank: ”Em yêu màu … Đồng bằng, rừng núi Biển đầy cá tôm, Bầu trời cao vợi”). A. Xanh (Blue). B. Cam (Orange). C. Trắng (White). D. Đỏ (Red). |
@@
| 1 | Wh-Questions | 445 |
| 2 | Questions with listed options | 96 |
| 3 | Yes/No questions | 5 |
| 4 | Others | 18 |
@@ 6.3 Discussion
| 1 | Passage: passage#01 Question: Trong ba đứa cháu, ai là đứa cháu được ông nhận xét là còn thơ dại? (In three children, who is claimed as innocent?) A. Vân B. Việt C. Xuân D. Cả 3 bạn trên (All of those above) | A | D |
| 2 | Passage: passage#02 Question: Dựa vào diễn biến sự việc và tâm trạng nhân vật, nên chia đoạn trích thành mấy đoạn nhỏ là hợp lí? (Based on the happenings of the event and the character’s mood, how many parts should the excerpt be divided into?) A. 2 đoạn (Trước cảnh thề nguyền – Cảnh thề nguyền) (2 parts (Before the oath - The oath)) B. 3 đoạn (Kiều sang nhà Kim Trọng – Kim tỉnh giấc – Thề nguyền) (3 parts (Kieu goes to Kim Trong’s house - Kim wakes up - The oath)) C. 4 đoạn (cứ 5 dòng thơ một đoạn) (4 parts (1 every 5 lines)) D. A và B đều được (A and B are correct) | A | D |
| 3 | Passage: passage#03 Question: Dấu câu nào không dùng để tách thành phần câu được mở rộng trong bài văn này? (What punctuation is not used to separate the additional information from sentences in this text?) A. Dấu ngoặc đơn (brackets) B. Dấu hai chấm (colon) C. Dấu phẩy (comma) D. Dấu ngoặc đơn và dấu phẩy (brackets and comma) | B | D |
| 4 | Passage: passage#04 Question:: Hãy sắp xếp các chi tiết dưới đây theo đúng thứ tự xuất hiện trong truyện Sơn Tinh, Thủy Tinh. (Let arrange these facts from the Son Tinh Thuy Tinh story by order of appearance) 1. Hùng Vương thứ mười tám nêu ra yêu cầu về lễ vật. (Hùng King XVIII gives the requirements of the marriage gift) 2. Sơn Tinh đem lễ vật đến trước và cưới được vợ. (Son Tinh comes first and married the bride) 3. Vua Hùng tổ chức kén rể cho Mị Nương. (Hùng King XVIII organizes a wedding for his daughter) 4. Sơn Tinh – Thủy Tinh đánh nhau ròng rã mấy tháng trời. (Son Tinh and Thuy Tinh keep fighting for months) A. (1) - (2) - (3) - (4). B. (1) - (3) - (2) - (4). C. (3) - (1) - (2) - (4). D. (1) - (3) - (4) - (2). | C | D |